OpenStudy (anonymous):

If 27 people can consume 54 kg of sweet potatoes within 45 days, in how many days does 72 people can consume 240 kg of sweet potatoes cosidering that they have the same rate of consuming?

OpenStudy (anonymous):

do you have the selections of answer?

OpenStudy (anonymous):

none...

OpenStudy (anonymous):

Use ratios :D 72/27 = 8/3 8/3 x 27 = 72 people 8/3 x 54 = 144 kg 8/3 x 45 = 120 days which gives us 72 ppl, eat 144 kg in 120 days use ratios again to adjust weight.. 240/144 = 5/3 5/3 x 144 = 240 kg 5/3 x 120 = 200 days therefore 72 ppl eat 240kg in 200days

OpenStudy (anonymous):

54/27=2 This is the kg of sweet potatoes 1 person eats in 45 days. 2/45=0.4444 This is the kg of sweet potatoes 1 person eats in 1 day. 72 x 0.4444 = 3.1999 The kg of sweet potatoes 72 people eat in 1 day. 3.1999(y) =240 y=240/3.1999 y=75 days

OpenStudy (anonymous):

@ Diane: 27 people eat 54 kg in 45 days; if we do a rough estimate and triple everything: 81 people eat 162 kg in 135 days.. even though they eat less the time required is much higher than 75days.

OpenStudy (anonymous):

cant find any problems with your reasoning though o.o except that it's 0.04444 not 0.4444

OpenStudy (anonymous):

81 people would eat 162 kg in 45 days.

OpenStudy (anonymous):

You're right it should be 0.0444 -- I mistyped. Thanks.

OpenStudy (anonymous):

Oh right i see, then: 81 would eat 162 in 45 days x1.5 81 would eat 243 in 67.5 days yup, my bad.